### v1

# The strong Fe K line and spin of the black-hole X-ray binary MAXI J1631–479

Abstract
We study the transient black hole binary MAXI J1631–479 in its soft spectral state observed simultaneously by the NICER and NuSTAR instruments. Its puzzling feature is the presence of a strong and broad Fe K line, while the continuum consists of a strong disk blackbody and a very weak power-law tail. The irradiation of the disk by a power-law spectrum fitting the tail is much too weak to account for the strong line. Two solutions were proposed in the past. One invoked an intrinsic Fe K disk emission, and the other invoked disk irradiation by the returning blackbody emission. We instead find that the strong line is naturally explained by the irradiation of the disk by the spectrum from Comptonization of the disk blackbody by coronal relativistic electrons. The shape of the irradiating spectrum at $\lesssim$10 keV reflects that of the disk blackbody; it is strongly curved and has a higher flux than that of a fit with a power-law irradiation. That flux accounts for the line. While this result is independent of the physical model used for the disk intrinsic emission, the value of the fitted spin strongly depends on it. When using a Kerr disk model for a thin disk with a color correction, the fitted spin corresponds to a retrograde disk, unlikely for a Roche-lobe overflow binary. Then, a model accounting for both the disk finite thickness and radiative transfer yields a spin of $a_{*}\approx 0.8$–0.9, which underlines the strong model-dependence of X-ray spin measurements.
show]aaz@camk.edu.pl
1 Introduction
[…]
Its BH mass, $M$, and the distance, $D$, remain uncertain. Rout et al. (2023) claimed a very high BH mass, $26{M}_{\sun}$, with a preference for $\sim\!60{M}_{\sun}$. This followed from a disk blackbody model at a high inclination, $i\approx 71\pm 1\arcdeg$, found by them. On the other hand, low inclination values were found in the spectral studies by Xu et al. (2020), $i=29\pm 3\arcdeg$, and by Draghis et al. (2024), $i=22^{+10}_{-12}\arcdeg$. Its dimensionless spin parameter was estimated to be very high, $a_{*}>0.94$ by Xu et al. (2020), and $a_{*}>0.996$ by Rout et al. (2023). Both studies used the X-ray reflection method (Bambi et al., 2021) in the soft spectral states of MAXI J1631–479. However, the results of this method crucially depend on the shape of the irradiating continuum, which is likely to be different from the usually assumed power law shape (Zdziarski et al., 2025a), especially in the soft state, when it is apparently formed by Compton upscattering of the disk emission by energetic electrons.
Here, we measure the spin using the continuum method (McClintock et al., 2014) coupled with the reflection method. This is applicable in the soft spectral state. The continuum method relies on the dependence of the innermost stable circular orbit (ISCO) radius on spin (Bardeen et al., 1972), utilizing relativistic models of accretion disks to fit observed X-ray spectra. For high-quality X-ray data, the coupled method can also be used to estimate the mass and distance (Parker et al., 2016), which we follow here.
We study a set of fully simultaneous NICER and NuSTAR observations of this source in the very soft state, i.e., ones with a strong disk blackbody and a very weak high-energy tail. We use two models of relativistic disks, kerrbb (Li et al., 2005) and slimbh (Straub et al., 2011). The former treats modifications of the local disk spectra with respect to the LTE using a color correction, defined as the ratio of the color temperature to the effective one. The latter uses calculations of the vertical radiative transfer of the disk emission (Davis et al., 2005; Davis & Hubeny, 2006), as well as taking into account the increasing disk scale height when the luminosity becomes comparable to the Eddington limit (e.g., Sądowski et al. 2011).
A puzzling feature of this source is the presence of a strong and broad Fe K emission line feature in the presence of a weak high-energy tail, especially in the initial part of the studied observation (Xu et al., 2020). When the tail is fitted by a power law plus reflection, the latter is way too weak to account for that feature. The initial solution proposed by Xu et al. (2020) was that the broad line is intrinsically emitted by the disk, being formed due to the transfer of the blackbody emission from deep in the disk to its surface. However, this disagrees with the standard disk spectra taking into account vertical radiative transfer (Davis et al. 2005; Davis & Hubeny 2006; S. Davis, private communication), which do not show strong Fe K features. Then, Rout et al. (2023) proposed that the Fe K feature is due to the reflection of returning blackbody radiation emitted by the disk (Schnittman & Krolik, 2009). However, Rout et al. (2023) modeled the returning radiation as a single blackbody with a free normalization (García et al., 2022), which can lead to spurious results. Furthermore, they included both the single blackbody and the disk blackbody in their fits, which is clearly incorrect. In the present work, we attempt to resolve this puzzle.
2 The data
The MAXI (Matsuoka et al., 2009) and Swift/BAT (Barthelmy et al., 2005) light curves of the outburst of MAXI J1631–479 are shown in Figure 1 of Xu et al. (2020). We study here the NuSTAR observation denoted as OBS1 in that paper (ObsID 90501301001), which was performed on MJD 58500, after the peak of the MAXI light curve and before the peak of the BAT light curve. Its spectra consist of a strong disk blackbody component and a weak high-energy tail. We define here Part I of the OBS1 as in Figure 1 and Table 1 of Xu et al. (2020). However, we define here Part II excluding a transitional period. In addition, we consider simultaneous observations during the same epochs by NICER, presented in Rout et al. (2021). The log of the observations studied here is given in Table 1, and the light curves of the detectors and the NuSTAR hardness ratio are shown in Figure 1.
The NuSTAR data are used in the 3–79 keV range. They were extracted using the option for bright sources222statusexpr=
"(STATUS==b0000xxx00xxxx000)&&(SHIELD==0)"
. Despite the NuSTAR and NICER being nearly simultaneous, we find the average spectrum from NICER shows some disagreement with the NuSTAR average spectrum in the overlapping energy range (see Section 3). In addition, the NICER spectrum shows a strong positive feature between 1 and 2 keV, which appears to be due to a calibration issue (see, e.g., Hall et al. 2025). Thus, we use the NICER spectrum in the energy range of 2–10 keV. The spectral data have been optimally binned (Kaastra & Bleeker, 2016) with an additional requirement of at least 25 counts per bin.
Figure 1: The NICER 2–10 keV and NuSTAR 3–10 and 10–60 keV light curves, and the NuSTAR hardness ratio during the studied observations. The zero time corresponds to the start of the NuSTAR observation, and the vertical dashed lines show the end of Part I (red) and the beginning of Part II (black; see Table 1).
Table 1: The log of the observations.
NICER Start Time Exposure (s) NuSTAR Start Time Exposure (s) Exposure (s)
Obs. ID End Time Obs. ID End Time FPM A FPM B
1200500103 2019-01-17 04:37:08 3016 90501301001 2019-01-17 02:38:58 6670 7018
Part I 2019-01-17 12:28:51 Part I 2019-01-17 13:17:04
1200500103 2019-01-17 16:58:30 4870 90501301001 2019-01-17 15:55:37 7471 7949
Part II 2019-01-17 23:24:54 Part II 2019-01-18 05:23:26
3 Spectral fits
3.1 The method
We follow the self-consistent soft-state spectral modeling developed by Zdziarski et al. (2025a). The main new feature of this method is the use of convolution for both calculating the spectrum from Comptonization of the disk emission and calculating the reflection spectrum from the former spectrum irradiating the disk. This avoids the assumption of the incident spectra being power laws with either exponential or thermal-Compton cutoffs, inherent when relxill or relxillCp models (García & Kallman, 2010; García et al., 2018b; Dauser et al., 2016) are used.
[…]
As found for Cyg X-1 (Zdziarski et al., 2024b) and GX 339–4 (Zdziarski et al., 2025a), high-energy tails in the soft state require the presence of nonthermal tails beyond the Maxwellian electron distribution of the Comptonizing plasma. The tail used in comppsc is a power law in the electron momentum, and it is given by $(\gamma\beta)^{-p}$ (where $\beta$ is the dimensionless electron velocity) from $\gamma_{\rm min}$, at which the normalizations of the Maxwellian and of the power law are equal, up to $\gamma_{\rm max}$. In our fits, we obtain relatively large values of $p$, for which the fits are independent of $\gamma_{\rm max}$, which we thus set equal to 100.
We use two models of relativistic disks. One is kerrbb (Li et al., 2005), which models geometrically thin disks using the formalism of Novikov & Thorne (1973) and assumes a local spectrum to be a blackbody with a color correction, $f_{\rm col}$, which we assume to equal 1.7 (Davis & El-Abd, 2019). The other is slimbh (described in Straub et al. 2011), based on the calculations of Sądowski (2009, 2011) and Sądowski et al. (2011). It takes into account the finite disk thickness, important for $L\gtrsim 0.1L_{\rm E}$ (Abramowicz et al., 1988). It also treats the local spectra using the radiative-transfer calculations of Davis et al. (2005) and Davis & Hubeny (2006) (when using the option $f_{\rm col}=-1$).
Our model for spectral fitting in xspec (Arnaud, 1996) is
${\tt plabs*tbfeo*comppsc}\{{\tt disk}+$
${\tt relconv[xilconv(comppsc}({\tt disk}))]\},$ (1)
where disk = kerrbb or slimbh. The first term in the curly brackets gives the disk emission alone, and the second term gives the relativistically broadened reflection. Both are subject to Comptonization, given by the comppsc term outside the brackets. The spin parameters in disk and relconv are tied. Then, plabs multiplies the model spectra by $KE^{-\Delta\Gamma}$, where $K=1$ and $\Delta\Gamma=0$ for the NuSTAR A unit, while they are free for the other detectors. The interstellar absorption is modeled by tbfeo (Wilms et al., 2000), where we assume the abundances of Anders & Grevesse (1989), except that those of O and Fe are free (with the relative abundances, $Z$, constrained to 0.5–2.0). We define our photon energy grid by the command energies 0.01 1000 1000 log (necessary for convolution models). We note that increasing the sampling of the energy grid twice (energies 0.01 1000 2000 log) did not influence the fit or its parameters. We use 90% confidence uncertainties ($\Delta\chi^{2}\approx 2.71$; Lampton et al. 1976).
3.2 Individual fits
The spectra of Part I and Part II were initially fitted separately. We first used kerrbb. We obtain good fits, with $\chi_{\nu}^{2}$ values of 531/422 and 638/506, respectively. However, the fitted spins, masses, and distances strongly disagree. In particular, $a_{*}=-0.70_{-0.10}^{+0.32}$, for Part I, while it is positive for Part II, $a_{*}=0.67_{-0.53}^{+0.15}$.
The agreement of the spin parameters is substantially better while using slimbh and the obtained values are positive; $a_{*}=0.85_{-0.07}^{+0.13}$ and $0.81_{-0.10}^{+0.07}$ for Part I and II, respectively. Also, the fits are better, $\chi_{\nu}^{2}=517/422$ and 635/506, respectively. Still, the fitted spins, masses, inclinations, and distances are different. We thus opt for joint fitting as our final approach (Section 3.3).
Before that, we show the presence of Fe K residuals and reflection using the individual data. We fit the NuSTAR data alone with a model without reflection, i.e., with Equation 1 without the second term in the curly brackets. We obtain bad fits with $\chi^{2}_{\nu}=1078/297$ and 1012/390, respectively, with strong residuals showing the presence of strong reflection. The data-to-model ratios are shown in Figure 2. We see distinct Fe K features around 7 keV, edges at energies above it, and reflection humps at energies $\gtrsim$10 keV, confirming the result of Xu et al. (2020). However, we see a clear discrepancy between the line profiles from FPM A and B for both data sets, with the latter showing dips around 7 keV. The apparent dip in the NuSTAR FPM B profile was attributed to the presence of an absorption line by Rout et al. (2023). However, this issue is discussed in detail in Draghis et al. (2024), who explain it as due to the proximity of the Sun causing source tracking to be unstable. They attempted to improve the data extraction using a custom procedure. However, the marked difference between the FPM A and B profiles around 7 keV remained, as shown in their Figure 19. Here, we use both detectors, but test for the effect of the removal of the FMP B in Section 3.3 below.
Refer to captionRefer to caption
Figure 2: The data to model ratios for the model fitted with slimbh to the NuSTAR data alone and without the reflection component for Part I (left) and Part II (right). We see patterns characteristic of Compton reflection.
3.3 Joint fitting
Refer to captionRefer to caption
Figure 3: The NICER (green) and NuSTAR FPM A and B (black and red, respectively) unfolded spectra (top panels) and data-to-model ratios (bottom panels) for the joint fit using slimbh. The Part I and II results are shown in the left and right panels, respectively. The total model spectra and the unabsorbed ones are shown by the solid black and blue curves, respectively. The unscattered parts of the disk emission, scattering alone, and reflection are shown by the magenta, green, and red curves, respectively. Their sums equal the total unabsorbed spectra (in blue).
Table 2: The results of the joint spectral fitting with slimbh
Component Parameter Part I Part II
Disk $L_{\rm disk}/L_{\rm E}$ $0.27^{+0.02}_{-0.02}$ $0.26^{+0.02}_{-0.06}$
Compton $kT_{\rm e}$ [keV] $18.4^{+9}_{-4.5}$ $8.1^{+1.4}_{-1.1}$
$y$ $028^{+0.016}_{-0.008} 0.0$$046^{+0.004}_{-0.007}$
$p$ $3.8_{-0.2}^{+0.2}$ $4.4_{-0.1}^{+0.2}$
$\gamma_{\rm min}$ $29^{+0.07}_{-0.03} 1.1$$12^{+0.02}_{-0.02}$
$f_{\rm cov}$ $0.36^{+0.09}_{-0.13}$ $0.57^{+0.10}_{-0.10}$
[…]
$\log_{10}\xi$ $4.3_{-0.1}^{+0.2}$ $4.3_{-0.1}^{+0.1}$
$q$ $3.8^{+0.2}_{-0.2}$ $3.5^{+0.2}_{-0.1}$
[…]
$Z_{\rm O,ISM}$ $0.86^{+0.26}_{-0.20}$
$Z_{\rm Fe,ISM}$ $0.50^{+0.08}$
$M\,[{M}_{\sun}]$ $9.2_{-1.0}^{+1.3}$
DD\,[kpc] $4.9^{+0.5}_{-0.5}$
$a_{*}$ $0.84^{+0.03}_{-0.03}$
$i\,[\arcdeg]$ $33^{+2}_{-1}$
$Z_{\rm Fe,disk}$ $4.6_{-0.6}^{+1.2}$
$\gamma_{\rm max}$ 100100f
$\chi_{\nu}^{2}$ 1169/938
Here, we fit the two data sets jointly. We assume common values of the ISM absorption parameters, mass, distance, inclination, spin, and the Fe abundance of the reflecting medium. We first use kerrbb. We obtain $\chi_{\nu}^{2}=1192/938$ giving a negative spin of $a_{*}=-0.61^{+0.23}_{-0.26}$. However, retrograde accretion is unlikely for an LMXB, see also the discussion of past claims of negative spins in LMXBs in Zdziarski et al. (2025a).
Given that, we decide on slimbh to be our final model, which also yields a better fit with $\chi_{\nu}^{2}=1169/938$, giving a relatively high and positive spin of $a_{*}=0.84\pm 0.03$; see Table 2 for all of the parameters. The corona covers roughly half of the disk, with $f_{\rm cov}\sim 0.5$. The obtained mass is within the typical range for BHs in LMXBs (Özel et al., 2010), and the inclination agrees with that of Xu et al. (2020). We note that slimbh does not allow for $a_{*}<0$. However, the $\chi^{2}$ dependence on the decreasing $a_{*}$ is approximately monotonically increasing, with $\Delta\chi^{2}=+34$ at $a_{*}=0$. Thus, the presence of a local minimum at $a_{*}<0$ appears unlikely. The spectra and residuals are shown in Figure 3. An important feature of these spectra is that the scattered disk emission, which is both emitted outside and impinges on the disk, has a shape very different from a power law. We see no significant residuals for the NuSTAR data (except for the highest energies), but we see that the NICER spectrum shows wavy residuals, indicative of instrumental issues. The apparent line above 2 keV is due to the gold edge of the NICER detector. Since its contribution to $\chi^{2}$ is small, we opted not to model it.
Then, we study the effect of the removal of the FMP B data (which can be affected by the proximity of the Sun, see Section 2). We found its effect on the fitting parameters to be minor, with most of the best-fit parameters having almost the same values, while the uncertainty ranges increase. The spin parameter becomes $a_{*}=0.83^{+0.05}_{-0.08}$. The largest difference is in the value of the reflection strength for Part I, which is ${\cal R}=1.8{+1.2}_{-0.7}$, extending to lower values than the full data, ${\cal R}=2.4^{+0.6}_{-0.4}$.
Finally, we consider a model with a warm and optically thick corona on top of the accretion disk, considered before for Cyg X-1, LMC X-1, M33 X-7, and GX 339–4 (Belczyński et al., 2024; Zdziarski et al., 2024a, b, 2025a). Such coronae were modeled there by optically thick thermal Comptonization using the code thcomp (Zdziarski et al., 2020). The differences of this model with respect to that of Equation 1 is the addition of the thcomp factor in front of the disk factor and that a free color correction in slimbh is used instead of the model including radiative transfer (see a discussion in Zdziarski et al. 2025a). However, we have not been able to find a model with a warm corona that would provide a better fit than our main model. The best one had $\chi_{\nu}^{2}=1180/933$. Thus, we conclude that such a component may not be present in MAXI J1631–479.
4 Discussion
Figure 4: The hybrid electron distributions corresponding to the joint best fit using comppsc and slimbh, see Figure 3 and Table 2. The distributions for Part I and II are shown by the red and blue solid curves, respectively. The dashed cyan and magenta curves show the corresponding Maxwellian spectra. The normalization is arbitrary.
We have obtained a relatively large Eddington ratio, $L_{\rm disk}/L_{\rm E}\approx 0.3$ (see Table 2), confirming the need to use a model of slim disks. The observed bolometric fluxes are 4.1 and $4.5\times 10^{-8}$ erg cm-2 s-1, respectively. However, the source is strongly absorbed, and the unabsorbed bolometric flux is about four times larger, $F_{\rm bol}=1.7$ and $1.8\times 10^{-7}$ erg cm-2 s-1, respectively. The flux in the disk alone (before scattering) is $\approx\!1.4\times 10^{-7}$ erg cm-2 s-1. At the best-fit of the distance and the inclination, the total unabsorbed fluxes correspond to the luminosity of $\approx\!3\times 10^{38}$ erg s-1, assuming the cosine dependence of the flux ($L=2\pi D^{2}F_{\rm bol}/\cos i$).
The reflection fractions are $\approx$2.4 and $\approx$0.8 for Parts I and II at the best fit, respectively, though the former can be significantly lower, ${\cal R}\approx 1.8{+1.2}_{-0.7}$, if the FMP B data are not used. The relatively high values for Part I can be explained by the effect of anisotropy of a slab corona containing relativistic electrons (Ghisellini et al., 1991). The seed blackbody photons from the disk are emitted upward and are scattered by the electrons in the corona. The highest probability of a collision is when electrons are moving downward. When the electrons are relativistic, they scatter preferentially forward, thus in the direction of the disk. This leads to enhancement of the reflection when measured with respect to the radiation emitted to the observer. The enhancement can be large, even by a factor of several (Ghisellini et al., 1991). The higher value for Part I is likely due to its coronal electrons being more relativistic than those of Part II. This is illustrated in Figure 4, which shows the fitted hybrid (Maxwellian followed by a power law) distributions. The hybrid electrons in our calculations have distinct high-energy tails, and a part of them is relativistic, especially for Part I with a large fraction of electrons with $\beta\gamma\gtrsim 1$, as shown in Figure 4. Thus, the anisotropy effect is strong and can easily account for the large reflection fraction (especially when the uncertainty related to the FMP B accuracy is taken into account). Furthermore, the corona for Part I has a much lower optical depth, with $\tau_{\rm T}\approx 0.19$, than $\tau_{\rm T}\approx 0.73$ for Part II. The scattering in Part I is thus mostly single, for which the relativistic kinematics forces it to be largely towards the disk, while $\tau_{\rm T}\sim 1$ for Part II, when multiple scatterings isotropise the emitted flux. Ideally, this effect could be strictly included in comppsc for the slab geometry, in which we could switch between the down and up emissions. A development of such an extension of comppsc is, however, beyond the scope of this work.
Our obtained value of the spin is relatively high, with the fitted $a_{*}\approx 0.8$–0.9. This is much more than most of the spins measured via gravitational waves emitted during mergers of two BHs in binaries, which are $a_{*}\approx 0.1\pm 0.1$ on average (The LIGO Scientific Collaboration et al., 2025). Also, it is more than the typical natal BH spin of $a_{*}\lesssim 0.1$ (Fuller & Ma, 2019; Belczyński et al., 2020). An increase of the spin from $a_{*}=0.1$ to $a_{*}=0.8$ requires past accretion of about 30% of the current BH mass, i.e., $\approx 3{M}_{\sun}$ (using the formulae of Bardeen 1970). This can be achieved only in binary evolution models with the initial masses of donors in BH LMXBs significantly higher than those presently observed (Podsiadlowski et al., 2003; Fragos & McClintock, 2015).
In this work, we modeled both the Comptonization and reflection self-consistently by using convolution models. We assumed that the high-energy tail is from Comptonization of the disk blackbody photons, which appears very likely given the strength of the disk blackbody component. Consequently, the scattered spectrum is strongly curved, with the shape far from a power law, see the green curve in Figure 3. The flux of this spectrum at the Fe K region is a factor of several times higher than the flux of the power-law fitted with reflection to the high-energy tail. Thus, the irradiating flux is enhanced by at least that factor, which, as we found, can explain the presence of the relatively strong Fe K line in the observed spectrum. This clearly demonstrates that spectral fits assuming the incident spectrum being a power law, such as relxill or reflionx, would give incorrect results for this source, as well as for other sources with soft-state high-energy tails. Instead, it is necessary to use convolution models for both Comptonization and reflection.
We are aware of the limitations of our treatment of reflection. The xilconv model is based on the tables used by xillver for the density of $10^{15}$ cm-3, while the disk surface density is likely to be much larger. This can result in an overestimated Fe abundance (García et al., 2018a), as in our case. The limited accuracy of xilconv was pointed out by Ding et al. (2024). Furthermore, the xilver tables have been done for a cold slab heated only by the irradiating photons, whereas in the present case the intrinsic dissipation dominates over heating by irradiation, and the maximum disk blackbody temperature is $kT_{\rm max}\approx 1$ keV (as we have checked replacing slimbh by diskbb (Mitsuda et al., 1984)). However, a more accurate convolution reflection model is not available at this time. Likely, it would have given different values of the fitted ionization parameter, the relative Fe abundance, and the spin. However, we are convinced that our main result, that the observed Fe K line can be explained by the irradiation by the spectrum from Compton scattering of disk blackbody photons, would stand.
Another caveat for our results is that we use the accretion disk model in the region of the Eddington ratio in which the disk is unstable (Lightman & Eardley, 1974; Shakura & Sunyaev, 1976). Accretion disks can be stabilized by large-scale magnetic fields, e.g., Sądowski (2016), Mishra et al. (2020). However, the structure of such disks can be quite different from the standard disks, e.g., Lančová et al. (2019). In particular the ISCO radius may not form a boundary for the flow, e.g., Rule et al. (2025). Still, the constancy of the inner radius in the soft state of BH XRBs found observationally in many sources in the soft state (Gierliński & Done, 2004) provides an argument for an important role played by the ISCO radius.
Our results also confirm the model dependence of the fitted spins of BH XRBs using different methods, see Zdziarski et al. (2025a, b). Here, we obtained $a_{*}=-0.61_{-0.26}^{+0.23}$ using kerrbb and $a_{*}=0.84\pm 0.03$ using slimbh, while both results are different from the two previous measurements giving $a_{*}>0.94$ using the reflection method alone. While we argue that the result with slimbh provides the most accurate results on the grounds of the accuracy of the disk model and the physical plausibility, we are aware of the caveats for it (as discussed above). An independent determination of the mass and the distance of MAXI J1631–479 is highly desirable for testing our model, though it is very difficult given the very high extinction toward the source (Section 1).
5 Conclusions
We have successfully modeled the strong and broad Fe K line in MAXI J1631–479 by reflection of the irradiation of the accretion disk by the spectrum from Comptonization of the disk blackbody by coronal relativistic electrons. We have taken into account the detailed shape of this spectrum, including its curvature reflecting the shape of the disk blackbody, see Figure 3. That flux is significantly higher than that corresponding to a fit with a power-law irradiation and reflection, e.g., using the reflection models relxillCp or reflionx. We have used convolution models for both Comptonization and reflection, and we stress the necessity to use such models in the soft state of BH XRBs.
The distribution of the coronal electrons is hybrid, with a Maxwellian followed by a power-law tail. In Part I of the observation, a significant part of the coronal electrons was relativistic, which led to an enhancement of the flux directed back to the disk with respect to that directed outward (Ghisellini et al., 1991).
In our modeling, we have used two relativistic disk models to measure the BH spin using the continuum fitting method and coupled that with the reflection method. Using the standard continuum model kerrbb, we have obtained a negative spin, $a_{*}=-0.61^{+0.23}_{-0.26}$. Then, we used slimbh, which accounts for both a finite disk scale height at Eddington ratios comparable to unity and for the vertical radiative transfer of the disk emission. That model gave us $a_{*}=0.84^{+0.03}_{-0.03}$, which we consider more likely (though still a significant systematic error) than the former. However, this underlines the strong model dependence inherent to fitting spins of BH XRBs, and we consider that value as tentative. With the slimbh model, we estimated the BH mass as $9.2_{-1.0}^{+1.3}{M}_{\sun}$, the distance as $4.9^{+0.5}_{-0.5}$ kpc, and the disk inclination of $33^{+2}_{-1}\arcdeg$. The source Eddington ratio is $\sim$0.3.
Our preferred spin value may result from accretion of a relatively large mass, $\sim\!3{M}_{\sun}$, onto the BH during stellar evolution, which could be explained by the model with initial mass masses of donors in LMXBs significantly higher than those observed (Podsiadlowski et al., 2003).
Acknowledgements
We thank Jorge Casares for a discussion on the donor of MAXI J1631–479. We acknowledge support from the Polish National Science Center grants 2019/35/B/ST9/03944 and 2023/48/Q/ST9/00138. MS acknowledges support from the National Science Center grant 2023/50/A/ST9/00527. BY is supported by the Natural Science Foundation of China (NSFC) grants 12322307, 12361131579, and 12273026, and by the Xiaomi Foundation/Xiaomi Young Talents Program. MP acknowledges support from the JSPS Postdoctoral Fellowship for Research in Japan, grant number P24712, as well as the JSPS Grants-in-Aid for Scientific Research-KAKENHI, grant number J24KF0244.